OpenStudy (anonymous):

If sin a = .25 and cos b = .25, which of the following is sin b/2 + cos a/2 A. 1.60 B. 1.32 C. .26 D. 1.06

jimthompson5910 (jim_thompson5910):

sin(a) = 0.25 sin^2(a) = (0.25)^2 ... square both sides sin^2(a) = 0.0625 ----------------------------------------- sin^2(a) + cos^2(a) = 1 0.0625 + cos^2(a) = 1 cos^2(a) = 1 - 0.0625 cos^2(a) = 0.9375 sqrt( cos^2(a) ) = sqrt(0.9375) cos(a) = 0.9682458

jimthompson5910 (jim_thompson5910):

so if sin(a) = 0.25, then cos(a) = 0.9682458 (approximately)

jimthompson5910 (jim_thompson5910):

use this idea to find sin(b)

OpenStudy (anonymous):

what would the equation look like? im confused on how to set it up.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

cos(b) = 0.25 cos^2(b) = 0.0625 then you plug it into sin^2(b) + cos^2(b) = 1 and solve for sin(b)

OpenStudy (anonymous):

Im still confused. Sorry i suck at math @jim_thompson5910

jimthompson5910 (jim_thompson5910):

basically the same steps will be followed, but instead of 'a' you use 'b'

jimthompson5910 (jim_thompson5910):

sin^2(b) + cos^2(b) = 1 sin^2(b) + 0.0625 = 1 sin^2(b) = 1-0.0625 sin^2(b) = 0.9375 sin(b) = sqrt(0.9375) sin(b) = 0.9682458

OpenStudy (anonymous):

so then I take sin b/2 and cos a/2 and add them? @jim_thompson5910

OpenStudy (anonymous):

it'd look like this, sin(0.9682458/2) + cos(0.9682458/2) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

now you would use the identity sin(b/2) = sqrt( (1-cos(b))/2 ) so, sin(b/2) = sqrt( (1-cos(b))/2 ) sin(b/2) = sqrt( (1-0.25)/2 ) sin(b/2) = sqrt( 0.75/2 ) sin(b/2) = sqrt( 0.375 ) sin(b/2) = 0.6123724

OpenStudy (anonymous):

okay im catching on @jim_thompson5910

jimthompson5910 (jim_thompson5910):

cos(a/2) = sqrt( (1+cos(a))/2 ) cos(a/2) = sqrt( (1+0.9682458)/2 ) cos(a/2) = sqrt( (1.9682458)/2 ) cos(a/2) = sqrt( 0.9841229 ) cos(a/2) = 0.992029687

OpenStudy (anonymous):

0.9682458+0.6123724 its gonna be A @jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct